pH and pOH calculations for weak acids and weak bases
1. What is the pH of a 0.55 M solution of trimethylamine (pKb= 4.13)
2. What is the pH of a 0.25 M solution of jaspersamine? The pKb value for jaspersamine 4.40.
3. What is the pH of a 0.15 M solution of a weak acid ammonium bromide given that the Kb value for ammonia is 1.8x10^-5.
Answer #1

Following is the - complete Answer -&- Explanation: for Question - 1 and Question - 2, of the given: Question Set:... in....typed format...

-------------------------------------------------------------------------

\RightarrowQuestion - 1:

\RightarrowAnswer:

For tri-methyl-amine:   pH = 11.805

\RightarrowExplanation:

Following is the complete Explanation: for the above: Answer.

  • Given:
  1. Molarity of Tri-methyl amine: Ma = 0.55 M ( mol/L )
  2. For  Tri methyl amine: value of pKb = 4.13
  • ​​​​​​​Step - 1:

​​​​​​​\Rightarrow  pKb = 4.13

\RightarrowKb = 10 -4.13 = 7.413 x 10-5   

  • Step - 2:

​​​​​​​\RightarrowWe know: Kb = [X+ ] x [OH - ] / [ tri-methyl amine ] = 7.413 x 10-5   

\Rightarrow [OH - ] = \sqrt{[Kb * [trimethyl amine]]} = ( \sqrt{[(7.413 * 10-5 ) x ( 0.55 ) ]} ) =  0.00638 M ( mol/L )

  • Step - 3:

​​​​​​​\RightarrowpOH = - log 10 [OH- ] = - log10 [ 0.00638 ] = 2.195

\Rightarrow pH =  ( 14.0 - pOH ) = ( 14.0 - 2.195 ) = 11.805

--------------------------------------------------------------------------------

\RightarrowQuestion - 2:

\RightarrowAnswer:

For jaspersamine:   pH = 11.5

\RightarrowExplanation:

Following is the complete Explanation: for the above: Answer.

  • Given:
  1. ​​​​​​​Molarity of jaspersamine: Ma = 0.25 M ( mol/L )
  2. For  jaspersamine: value of pKb = 4.40
  • ​​​​​​​Step - 1:

​​​​​​​\Rightarrow  pKb = 4.40

\RightarrowKb = 10 -4.40 = 3.98 x 10 -5   

  • Step - 2:

​​​​​​​\RightarrowWe know: Kb = [X+ ] x [OH - ] / [ jaspersamine ] = 3.98 x 10 -5

\Rightarrow [OH - ] = \sqrt{[Kb * [jaspersamine]]} = ( \sqrt{[(3.98 * 10^{-5} ) * ( 0.25 ) ]} ) =  0.00315 M ( mol/L )

  • Step - 3:

​​​​​​​\RightarrowpOH = - log 10 [OH- ] = - log10 [ 0.00315 ] = 2.50

\Rightarrow pH =  ( 14.0 - pOH ) = ( 14.0 - 2.50 ) = 11.5  

-------------------------------------------------------------------------------

Note: The remaining question: i.e. Question - 3: can be solved, by applying the method followed for solving the above stated solutions...

​​​​​​​